OpenStudy (anonymous):

Graph the line with slope -3 passing through the point (3,3)

OpenStudy (anonymous):



OpenStudy (anonymous):

(y - y1) = m(x - x1) y - 3 = -3(x -3) y - 3 = -3x + 9 y = -3x + 12 put x = 0 then y = 12 so (0, 12) put y = 0 then x = 4 so (4, 0) plot these points n u hv the line passing through (3,3) and having slope -3

OpenStudy (anonymous):

see attached..

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

U r welcome. U could show yr appreciation by awarding a medal to the people whose answer you found satisfactory. That will also mark the question as closed/answered satisfactorily....☺